OpenStudy (bruno102):

How would I attempt to solve this? You have a pool that measure 15 yards by 10 yards. You want to build a deck around all four sides of the pool and you want to increase the total area to 250 square yards. What is the width of the deck?

OpenStudy (anonymous):

Let w = width of deck Then area of pool + area of deck = 250 yd² (15 + 2w) * (10 + 2w) = 250 4w² + 50w + 150 = 250 4w² + 50w - 100 = 0 w = (-50 ± √(2500+1600)) / 8 -----> reject negative solution w = (-50 + 10√41) / 8 w = 10 (-5 + √41) / 8 w = 1.25 (-5 + √41) w = 1.753905297 Width of deck is approximately 1.75 yds